OpenStudy (anonymous):

-4 sqrt x+2+3= -1 Solve for x???

OpenStudy (lgbasallote):

\[-4\sqrt{x+2} + 3 = -1\] subtract 3 ffrom both sides \[-4\sqrt{x+2} = -1 - 3\] \[-4 \sqrt{x-2} = -4\] divide bothsides by -4 \[\sqrt{x-2} = 1\] square both sides \[x- 2 = 1\] add 2 to both sides \[x = 1 + 2\] \[x = 3\] do you get that?

OpenStudy (anonymous):

but my options are : x=-2 x=-3 x=-2 and x=-3 None of the above

OpenStudy (anonymous):

??

OpenStudy (lgbasallote):

uhh i saw what i did wrong...my typo misled me

OpenStudy (lgbasallote):

\[-4\sqrt{x+2} + 3 = -1\] subtract 3 from both sides \[-4\sqrt{x+2} = -1 - 3\] \[-4\sqrt{x+2} = -4\] divide both sides by -4 \[\sqrt{x+2} = 1\] square both sides \[x + 2 = 1\] subtract 2 from both sides \[x = 1 - 2\] \[x = -1\]

OpenStudy (lgbasallote):

now do you get it?

OpenStudy (anonymous):

so it is none of the above? because there is no -1 for my options

OpenStudy (lgbasallote):

that is of course assuming the question is \[-4 \sqrt{x+2} + 3 = -1\]

OpenStudy (lgbasallote):

^is that the question?

OpenStudy (anonymous):

yes, that is the question, I worked it out just like you did and was all confused cause there wasn't a -1... agh so it is none of the above

OpenStudy (lgbasallote):

yup